An ACO is training registered nurses to conduct annual wellness visits, a service that is traditionally performed by physicians. At the current levels of nurses and physicians, if another nurse is hired for an hour, it will take nurses 30 minutes to perform an annual wellness visit, whereas if another physician is hired for an hour, it will take physicians 15 minutes to perform an annual wellness visit. At this ACO, all nurses are paid $50 per hour, and physicians are paid $100 per hour.

What is the marginal product (MP) for an additional hour of labor for nurses? The units are exams per hour.

What is the marginal product (MP) for an additional hour of labor for physicians? The units are exams per hour.

If the ACO were to dedicate additional funds to hiring either additional nurses or additional physicians, which type of health worker should it hire assuming its goal is productive efficiency?
Answer #1

a)

Each additional nurse can complete one wellness visit in 30 minutes i.e. nurse can compete 2 wellness visits in one hour.

So, Marginal Product of Nurse=MPN=2 exam/hour

b)

Each additional physician can complete one wellness visit in 15 minutes i.e. physician can compete 4 wellness visits in one hour.

So, Marginal Product of physician=MPP=4 exam/hour

c)

MPN/MPP=2/4=1/2

wN/wP=50/100=1/2

Since MPN/MPP=wN/wP, its a state of productive efficiency.

Firm can hire either of them.